Circular No. 6623

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997br IN ESO 576-40 The Beijing Astronomical Observatory (BAO) Supernova Survey (cf.

IAUC 6612) reports the discovery on Apr. 10.6 UT with the 0.6-m BAO reflector (+ unfiltered CCD) of a supernova (mag ∼ 14.5) located at α = 13^{h}20^{m}42^{s}.40, δ = −22^{o}02′12′′.3 (equinox 2000.0), which is 20′′.6 west and 51′′.6 north of the nucleus of ESO 576-40 = PGC 46574. Unfiltered CCD images of the same field taken on Jan. 8 (limiting mag ∼ 18.0) show no star at this position. Similar images taken on Apr. 11.6 show that SN 1997br had brightened to mag 14.2. A Low-resolution spectrogram (20 nm/mm, range 370–840 nm), obtained by Q.-y. Qiao, H. Wu, J.-y. Wei, and W.-d. Li (BAO) with the BAO 2.16-m telescope at Xinglong station, shows this to be a peculiar type-Ia supernova prior to optical maximum. The spectrum is rather blue, with the whole appearance similar to SN 1991T prior to optical maximum. There is only very weak Si II 635-nm absorption at 628 nm. Other features include Ca II, Mg II, Si II, and S II lines, with the minima measured at 396, 427, 497, and 572 nm, respectively. There is conspicuous host-galactic Na II absorption at 591.8 nm (equivalent width ∼ 0.067 nm), showing that the supernova is reddened by the host galaxy. The expansion velocity of the supernova envelope is measured to be ∼ 12 000 km/s, as derived from the Si II absorption at 497.2 nm. The redshift of the host galaxy is ∼ 2280 km/s, from the narrow Hα line and the Na II host-galactic absorption line.

SUPERNOVA 1997bq IN NGC 3147 P. Garnavich and R. Kirshner, Center for Astrophysics, report that

CCD images of SN 1997bq taken by N. Grogin and D. Moraru with the Whipple Observatory’s 1.2-m telescope between Apr. 8 and 11 show that it has brightened by 0.7 mag. This indicates the supernova is still well before maximum and may reach V = 14.5. The estimated magnitude on Apr. 11.3 UT was V = 15.3.

SUPERNOVA 1997bp in NGC 4680 R. Evans, Coonabarabran, N.S.W., reports a 0.2-mag visual bright-

ening of this supernova between Apr. 7.5 and 9.5 UT. Additional visual magnitude estimate by L. Kiss, Szeged, Hungary: Apr. 9.85 UT, 13.9.

1997 April 11  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green